### v1

@@ 5 Construction of Quantum Stabilizer Codes
Let $q=2$, $m=6$, $g_{1}(x)=x^{4}+x^{3}+x+1$, $g_{2}(x)=x^{5}+x^{4}+x^{3}+x^{2}+x+1$, $v_{1}(x)=x^{5}$, $v_{2}(x)=x^{5}+x$, we have $\gcd(v_{1}(x)v_{2}(x)-1,x^{m}-1)=1$. Let $\mathcal{C}$ be the QC code generated by $(g_{1}(x),v_{1}(x)g_{1}(x))$ and $(v_{2}(x)g_{2}(x),g_{2}(x))$. Through calculations in Magma, it is found that the parameters of linear code $\mathcal{C}^{\perp_{S}}$ are $[12,9,2]_{2}$. Applying Theorem 16, we construct a quantum error-correcting code with parameters $[[6,3,2]]_{2}$, which coincide with those of a best-known code listed in Grassl’s code table codetables .
@@ 6 Construction of Quantum Synchronizable Codes
duchao2023 Let $\mathcal{C}_{1}$ be a 2-generator QC code over $\mathbb{F}_{q}$ with parameters $[2m,k_{1},d_{1}]$, generated by $\{([a_{1,1}(x)],[a_{1,2}(x)]),([a_{2,1}(x)],[a_{2,2}(x)])\}$. Similarly, let $\mathcal{C}_{2}$ be another 2-generator QC code over $\mathbb{F}_{q}$ with parameters $[2m,k_{2},d_{2}]$, generated by $\{([b_{1,1}(x)],[b_{1,2}(x)]),([b_{2,1}(x)],[b_{2,2}(x)])\}$. Suppose that $\mathcal{C}_{1}^{\perp}\subset\mathcal{C}_{1}\subset\mathcal{C}_{2}$, and that for some integer $1\leq j\leq 2$, the greatest common divisor $\eta(x)=\gcd(a_{1,j}(x),a_{2,j}(x))$ is nontrivial, and satisfies $\eta(x)=f(x)b_{1,j}(x)$, where $b_{1,j}(x)=\gcd(b_{1,j}(x),b_{2,j}(x))$. Then, for any pair of non-negative integers $a_{l}$ and $a_{r}$ satisfying $a_{l}+a_{r}<\text{ord}(f(x))$, there exists an $(a_{l},a_{r})$-quantum synchronizable code with parameters $[[2m+a_{l}+a_{r},2(k_{1}-m)]]_{q}$, which can correct at least up to $\lfloor\frac{d_{1}-1}{2}\rfloor$ phase errors and at least up to $\lfloor\frac{d_{2}-1}{2}\rfloor$ bit errors.
When $\text{ord}(f(x))$ in the above lemma attains its maximum possible value $m$, the resulting quantum synchronizable code achieves the best possible tolerance against misalignment. We now focus on the cases where $m=p^{t}$ and $m=lp^{t}$, with $l,t\geq 1$ and $\gcd(l,p)=1$, providing explicit constructions for codes that attain this level of tolerance.
@@ 6.1 Quantum Synchronizable Codes from QC Codes with $m=p^{t}$
Let $\mathcal{C}_{1}$ be a QC code $[2p^{t},k_{1},d_{1}]$, generated by $([(x-1)^{r_{1}}],[v_{1}(x)(x-1)^{r_{1}}])$ and $([v_{2}(x)(x-1)^{r_{2}]},[(x-1)^{r_{2}}])$, where $v_{1}(x)$ and $v_{2}(x)$ are monic polynomials in $\mathcal{R}$ , $\gcd(v_{1}(x)v_{2}(x)-1,x^{p^{t}}-1)=1$, and $0<r_{1}<r_{2}<\frac{p^{t}-1}{2}$. Similarly, let $\mathcal{C}_{2}$ be another QC code $[2p^{t},k_{2},d_{2}]$, generated by $([(x-1)^{s_{1}}],[v_{1}(x)(x-1)^{s_{1}}])$ and $([v_{2}(x)(x-1)^{s_{2}}],[(x-1)^{s_{2}}])$, where $0<s_{1}<r_{1}$ and $0<s_{1}<s_{2}<r_{2}$. Suppose $r_{1}-s_{1}>p^{t-1}$. Then, for any non-negative integers $a_{l},a_{r}$ satisfying $a_{l}+a_{r}<p^{t}$, there exists a quantum synchronizable code with parameters
@@
Let $q=p=3$, $t=3$, $v_{1}(x)=x^{3}+x^{2}+1$, and $v_{2}(x)=x^{5}+2x^{3}+x^{2}+2x+1$. We choose $r_{1}=12$, $r_{2}=13$, $s_{1}=2$, and $s_{2}=5$. Then, $\mathcal{C}_{1}$ is generated by $\bigl{(}[(x-1)^{12}],[v_{1}(x)(x-1)^{12}]\bigr{)},\bigl{(}[v_{2}(x)(x-1)^{13}],[(x-1)^{13}]\bigr{)}$. Similarly, $\mathcal{C}_{2}$ is generated by $\bigl{(}[(x-1)^{2}],[v_{1}(x)(x-1)^{2}]\bigr{)}$, $\bigl{(}[v_{2}(x)(x-1)^{5}],[(x-1)^{5}]\bigr{)}$. By Theorem 17, we obtain a quantum synchronizable code with parameters $(a_{l},a_{r})-[[54+a_{l}+a_{r},4]]_{3}$, where $a_{l}+a_{r}<27$.
@@
Let $q=p=5$, $t=2$, $v_{1}(x)=x^{3}+x^{2}+2x+3$, and $v_{2}(x)=x^{5}+2x^{4}+x^{3}+2x^{2}+x$. We choose $r_{1}=9$, $r_{2}=11$, $s_{1}=3$, and $s_{2}=4$. Then, $\mathcal{C}_{1}$ is generated by $\bigl{(}[(x-1)^{9}],[v_{1}(x)(x-1)^{9}]\bigr{)},\bigl{(}[v_{2}(x)(x-1)^{11}],[(x-1)^{11}]\bigr{)}$. Similarly, $\mathcal{C}_{2}$ is generated by $\bigl{(}[(x-1)^{3}],[v_{1}(x)(x-1)^{3}]\bigr{)}$, $\bigl{(}[v_{2}(x)(x-1)^{4}],[(x-1)^{4}]\bigr{)}$. By Theorem 17, we obtain a quantum synchronizable code with parameters $(a_{l},a_{r})-[[50+a_{l}+a_{r},10]]_{5}$, where $a_{l}+a_{r}<25$.
@@ 6.2 Quantum Synchronizable Codes from QC Codes with $m=lp^{t}$
In this case, we set $m=lp^{t}$, where $\gcd(l,p)=1$. We begin by reviewing some basic concepts related to cyclotomic cosets Huffman2003 .
@@
$\bigl{(}[\prod_{s\in T_{l}}M_{s}(x)^{r_{1s}}],[v_{1}(x)\prod_{s\in T_{l}}M_{s}(x)^{r_{1s}}]\bigr{)},\quad\bigl{(}[v_{2}(x)\prod_{s\in T_{l}}M_{s}(x)^{r_{2s}}],[\prod_{s\in T_{l}}M_{s}(x)^{r_{2s}}]\bigr{)},$
@@
$\bigl{(}\prod_{s\in T_{l}}M_{s}(x)^{p^{t}-{r_{1(-s)}}},-$ $\overline{v}_{2}(x)\prod_{s\in T_{l}}M_{s}(x)^{p^{t}-{r_{1(-s)}}}\bigr{)},$
$\bigl{(}-\overline{v}_{1}(x)\prod_{s\in T_{l}}M_{s}(x)^{p^{t}-{r_{2(-s)}}},\prod_{s\in T_{l}}M_{s}(x)^{p^{t}-{r_{2(-s)}}}\bigr{)}.$
@@
Let $\mathcal{C}_{1}$ be the 2-generator QC code $[2lp^{t},k_{1},d_{1}]$, generated by
$\bigl{(}[\prod_{s\in T_{l}}M_{s}(x)^{r_{1s}}],[v_{1}(x)\prod_{s\in T_{l}}M_{s}(x)^{r_{1s}}]\bigr{)},\quad\bigl{(}[v_{2}(x)\prod_{s\in T_{l}}M_{s}(x)^{r_{2s}}],[\prod_{s\in T_{l}}M_{s}(x)^{r_{2s}}]\bigr{)},$
where $\gcd(v_{1}(x)v_{2}(x)-1,x^{lp^{t}}-1)=1$ and for any $s\in T_{l}$, we have $0<r_{1s}<r_{2s}<p^{t}-r_{2(-s)}$. Let $\mathcal{C}_{2}$ be the 2-generator QC code $[2lp^{t},k_{2},d_{2}]$, generated by
$\bigl{(}[\prod_{s\in T_{l}}M_{s}(x)^{j_{1s}}],[v_{1}(x)\prod_{s\in T_{l}}M_{s}(x)^{j_{1s}}]\bigr{)},\quad\bigl{(}[v_{2}(x)\prod_{s\in T_{l}}M_{s}(x)^{j_{2s}}],[\prod_{s\in T_{l}}M_{s}(x)^{j_{2s}}]\bigr{)},$
@@
Therefore, we obtain that $\mathcal{C}_{1}\subset\mathcal{C}_{2}$. In fact, it follows that $\eta(x)=\gcd\bigl{(}\prod_{s\in T_{l}}M_{s}(x)^{r_{1s}},v_{2}(x)\prod_{s\in T_{l}}M_{s}(x)^{r_{2s}}\bigr{)}=\prod_{s\in T_{l}}M_{s}(x)^{r_{1s}}$. Moreover,
@@ 7 Conclusion
In this paper, we first derived the necessary and sufficient conditions for the self-orthogonality of QC codes under Euclidean, symplectic, and Hermitian inner products. Building on this, we applied these conditions to a special class of 2-generator QC codes to characterize their self-orthogonal and dual-containing properties. Furthermore, we applied these results to construct quantum stabilizer codes and quantum synchronizable codes with good parameters. Future work may explore extensions to broader classes of QC codes and their applications in quantum error-correcting codes.
Acknowledgement
This work was supported by Shandong Provincial Natural Science Foundation of China (ZR2023LLZ013, ZR2022MA061), Fundamental Research Funds for Central Universities (23CX03003A) and China Education Innovation Research Fund (2022BL027).

### v2

@@ 5 Construction of quantum stabilizer codes
Let $q=2$, $m=6$, $g_{1}(x)=x^{4}+x^{3}+x+1$, $g_{2}(x)=x^{5}+x^{4}+x^{3}+x^{2}+x+1$, $v_{1}(x)=x^{5}$, $v_{2}(x)=x^{5}+x$, we have $\gcd(v_{1}(x)v_{2}(x)-1,x^{m}-1)=1$. Let $\mathcal{C}$ be the QC code generated by $(g_{1}(x),v_{1}(x)g_{1}(x))$ and $(v_{2}(x)g_{2}(x),g_{2}(x))$. Through calculations in Magma, it is found that the parameters of linear code $\mathcal{C}^{\perp_{S}}$ are $12,9,2]_{2}$. Applying Theorem 8, we construct a quantum error-correcting code with parameters $6,3,2]]_{2}$, which coincide with those of a best-known code listed in the codetable of Grassl [36].
@@ 6 Construction of quantum synchronizable codes
[34] Let $\mathcal{C}_{1}$ be a 2-generator QC code over $\mathbb{F}_{q}$ with parameters $2m,k_{1},d_{1}]$, generated by $\{([a_{1,1}(x)],[a_{1,2}(x)]),([a_{2,1}(x)],[a_{2,2}(x)])\}$. Similarly, let $\mathcal{C}_{2}$ be another 2-generator QC code over $\mathbb{F}_{q}$ with parameters $2m,k_{2},d_{2}]$, generated by $\{([b_{1,1}(x)],[b_{1,2}(x)]),([b_{2,1}(x)],[b_{2,2}(x)])\}$. Suppose that $\mathcal{C}_{1}^{\perp}\subset\mathcal{C}_{1}\subset\mathcal{C}_{2}$, and that for some integer $1\leq j\leq 2$, the greatest common divisor $\eta(x)=\gcd(a_{1,j}(x),a_{2,j}(x))$ is nontrivial, and satisfies $\eta(x)=f(x)b_{1,j}(x)$, where $b_{1,j}(x)=\gcd(b_{1,j}(x),b_{2,j}(x))$. Then, for any pair of non-negative integers $a_{l}$ and $a_{r}$ satisfying $a_{l}+a_{r}<\text{ord}(f(x))$, there exists an $(a_{l},a_{r})$-quantum synchronizable code with parameters $2m+a_{l}+a_{r},2(k_{1}-m)]]_{q}$, which can correct at least up to $\lfloor\frac{d_{1}-1}{2}\rfloor$ phase errors and at least up to $\lfloor\frac{d_{2}-1}{2}\rfloor$ bit errors.
Here, $\text{ord}(f(x))$ refers to the order of $f(x)$, namely the smallest positive integer $\tau$ such that $f(x)\mid x^{\tau}-1$ over the given field. When $\operatorname{ord}(f(x))$ in the above lemma attains its maximum possible value, the resulting quantum synchronizable code achieves the best possible tolerance against misalignment. We now focus on the cases where $m=p^{t}$ and $m=lp^{t}$, with $l,t\geq 1$ and $\gcd(l,p)=1$, providing explicit constructions for codes that attain this level of tolerance.
@@ 6.1 Quantum synchronizable codes from QC codes with $m=p^{t}$
Let $\mathcal{C}_{1}$ be a $2p^{t},k_{1},d_{1}]$ QC code generated by $([(x-1)^{r_{1}}],[v_{1}(x)(x-1)^{r_{1}}])$ and $([v_{2}(x)(x-1)^{r_{2}]},[(x-1)^{r_{2}}])$, where $v_{1}(x)$ and $v_{2}(x)$ are monic polynomials in $\mathcal{R}$ , $\gcd(v_{1}(x)v_{2}(x)-1,x^{p^{t}}-1)=1$, and $0<r_{1}<r_{2}<\frac{p^{t}-1}{2}$. Similarly, let $\mathcal{C}_{2}$ be another $2p^{t},k_{2},d_{2}]$ QC code generated by $([(x-1)^{s_{1}}],[v_{1}(x)(x-1)^{s_{1}}])$ and $([v_{2}(x)(x-1)^{s_{2}}],[(x-1)^{s_{2}}])$, where $0<s_{1}<r_{1}$ and $0<s_{1}<s_{2}<r_{2}$. Suppose $r_{1}-s_{1}>p^{t-1}$. Then, for any non-negative integers $a_{l},a_{r}$ satisfying $a_{l}+a_{r}<p^{t}$, there exists a quantum synchronizable code with parameters
@@
Let $q=p=3$, $t=3$, $v_{1}(x)=x^{3}+x^{2}+1$, and $v_{2}(x)=x^{5}+2x^{3}+x^{2}+2x+1$. We choose $r_{1}=12$, $r_{2}=13$, $s_{1}=2$, and $s_{2}=5$. Then, $\mathcal{C}_{1}$ is generated by $\bigl{(}[(x-1)^{12}],[v_{1}(x)(x-1)^{12}]\bigr{)}\enspace\text{and}\enspace\bigl{(}[v_{2}(x)(x-1)^{13}],[(x-1)^{13}]\bigr{)}$. Similarly, $\mathcal{C}_{2}$ is generated by $\bigl{(}[(x-1)^{2}],[v_{1}(x)(x-1)^{2}]\bigr{)}$ and $\bigl{(}[v_{2}(x)(x-1)^{5}],[(x-1)^{5}]\bigr{)}$. By Theorem 9, we obtain a quantum synchronizable code with parameters $(a_{l},a_{r})-[[54+a_{l}+a_{r},4]]_{3}$, where $a_{l}+a_{r}<27$.
@@
Let $q=p=5$, $t=2$, $v_{1}(x)=x^{3}+x^{2}+2x+3$, and $v_{2}(x)=x^{5}+2x^{4}+x^{3}+2x^{2}+x$. We choose $r_{1}=9$, $r_{2}=11$, $s_{1}=3$, and $s_{2}=4$. Then, $\mathcal{C}_{1}$ is generated by $\bigl{(}[(x-1)^{9}],[v_{1}(x)(x-1)^{9}]\bigr{)}$ and $\bigl{(}[v_{2}(x)(x-1)^{11}],[(x-1)^{11}]\bigr{)}$. Similarly, $\mathcal{C}_{2}$ is generated by $\bigl{(}[(x-1)^{3}],[v_{1}(x)(x-1)^{3}]\bigr{)}$ and $\bigl{(}[v_{2}(x)(x-1)^{4}],[(x-1)^{4}]\bigr{)}$. By Theorem 9, we obtain a quantum synchronizable code with parameters $(a_{l},a_{r})-[[50+a_{l}+a_{r},10]]_{5}$, where $a_{l}+a_{r}<25$.
@@ 6.2 Quantum synchronizable codes from QC codes with $m=lp^{t}$
In this case, we set $m=lp^{t}$, where $\gcd(l,p)=1$. We begin by reviewing some basic concepts related to cyclotomic cosets [38].
@@
$\bigl{(}[\prod_{s\in T_{l}}M_{s}(x)^{r_{1s}}],[v_{1}(x)\prod_{s\in T_{l}}M_{s}(x)^{r_{1s}}]\bigr{)}\enspace\text{and}\enspace\bigl{(}[v_{2}(x)\prod_{s\in T_{l}}M_{s}(x)^{r_{2s}}],[\prod_{s\in T_{l}}M_{s}(x)^{r_{2s}}]\bigr{)},$
@@
$\bigl{(}[\prod_{s\in T_{l}}M_{s}(x)^{p^{t}-{r_{1(-s)}}}],[$- $\overline{v}_{2}(x)\prod_{s\in T_{l}}M_{s}(x)^{p^{t}-{r_{1(-s)}}}]\bigr{)}\enspace\text{and}$
$\bigl{(}[-\overline{v}_{1}(x)\prod_{s\in T_{l}}M_{s}(x)^{p^{t}-{r_{2(-s)}}}],[\prod_{s\in T_{l}}M_{s}(x)^{p^{t}-{r_{2(-s)}}}]\bigr{)}.$
@@
Let $\mathcal{C}_{1}$ be the 2-generator $2lp^{t},k_{1},d_{1}]$ QC code generated by
$\bigl{(}[\prod_{s\in T_{l}}M_{s}(x)^{r_{1s}}],[v_{1}(x)\prod_{s\in T_{l}}M_{s}(x)^{r_{1s}}]\bigr{)}\enspace\text{and}\enspace\bigl{(}[v_{2}(x)\prod_{s\in T_{l}}M_{s}(x)^{r_{2s}}],[\prod_{s\in T_{l}}M_{s}(x)^{r_{2s}}]\bigr{)},$
where $\gcd(v_{1}(x)v_{2}(x)-1,x^{lp^{t}}-1)=1$ and for any $s\in T_{l}$, we have $0<r_{1s}<r_{2s}<p^{t}-r_{2(-s)}$. Let $\mathcal{C}_{2}$ be the 2-generator $2lp^{t},k_{2},d_{2}]$ QC code generated by
$\bigl{(}[\prod_{s\in T_{l}}M_{s}(x)^{j_{1s}}],[v_{1}(x)\prod_{s\in T_{l}}M_{s}(x)^{j_{1s}}]\bigr{)}\enspace\text{and}\enspace\bigl{(}[v_{2}(x)\prod_{s\in T_{l}}M_{s}(x)^{j_{2s}}],[\prod_{s\in T_{l}}M_{s}(x)^{j_{2s}}]\bigr{)},$
@@
Therefore, we obtain that $\mathcal{C}_{1}\subset\mathcal{C}_{2}$. In fact, it follows that
$\eta(x)=\gcd\bigl{(}\prod_{s\in T_{l}}M_{s}(x)^{r_{1s}},v_{2}(x)\prod_{s\in T_{l}}M_{s}(x)^{r_{2s}}\bigr{)}=\prod_{s\in T_{l}}M_{s}(x)^{r_{1s}}.$
Moreover,
@@ 7 Conclusion
In this paper, we first derived the necessary and sufficient conditions for the self-orthogonality of QC codes under Euclidean, symplectic, and Hermitian inner products. Building on this, we applied these conditions to a special class of 2-generator QC codes to characterize their self-orthogonal and dual-containing properties. Furthermore, we applied these results to construct quantum stabilizer codes and quantum synchronizable codes with good parameters. In future work, we plan to further investigate the conditions for quasi-twisted (QT) codes to be self-orthogonal via their generating sets. Moreover, we aim to apply the construction method proposed by Ezerman et al. in [32] to obtain quantum error-correcting codes with larger minimum distance.
[…]
Acknowledgements We would like to express our sincere gratitude to Professor Markus Grassl for his valuable comments and constructive suggestions, which have greatly contributed to improving the quality of this work.
Declarations
  • Funding This work was supported by Shandong Provincial Natural Science Foundation of China (ZR2023LLZ013, ZR2022MA061), Fundamental Research Funds for the Central Universities (23CX03003A), National Natural Science Foundation of China (61902429), and China Education Innovation Research Fund (2022BL027).
  • Conflict of interest The authors declare no conflict of interest.
  • Ethics approval and consent to participate Not applicable.
  • Consent for publication Not applicable.
  • Data availability Not applicable.
  • Materials availability Not applicable.
  • Code availability Not applicable.
  • Author contribution Authors Mengying Gao and Yuhua Sun were mainly responsible for the theoretical analysis and experimental design, while authors Tongjiang Yan and Chun’e Zhao contributed primarily to manuscript writing and table preparation.